Solve for X
4x to the 3rd -25x=0

Answer by checkley77(12844)

4x^3-25x=0
X(4X^2-25)=0
X(2X+5)(2X-5)=0
X=0 ANS.
2X+5=0
2X=-5
X=-5/2=-2.5 ANS.
2X-5=0
2X=5
X=5/2=2.5 ANS.